OpenStudy (habbster):

the dimensions of a rectangular solid are in the ratio of 1:6:6 and its volume is 562.5 cm^3 , find the measurement of the edges?

OpenStudy (phi):

cozy has the right idea, but volume is found by multiplying the 3 dimensions The 1:6:6 ratio means that the lengths are 1x , 6x and 6x the volume will be 36x^3= 562.5 cc solve for x

OpenStudy (habbster):

the 1:6:6 is the length the width and the height's ratio and the 562.5 is the volume so if you multiply the all the outcome of the ratio you should come up with 562.5

OpenStudy (habbster):

i come up with 15.625 but if i multiply it i wont give the volume that is given

OpenStudy (phi):

you found x^3 x^3= 5652.5/36 = 15.625

OpenStudy (habbster):

do i need to cube root it?

OpenStudy (phi):

A calculator gives the answer, but if you write it as an improper fraction \[ x^3 = \frac{125}{8} \]

OpenStudy (habbster):

yay thank you hahaha all i forgot is to cube root it

OpenStudy (phi):

to solve, raise both sides to the 1/3 power (that means cube root) \[ \left(x^3\right)^{\frac{1}{3}} = \left(\frac{125}{8}\right)^{\frac{1}{3}}\] \[ x= \frac{125^{\frac{1}{3}}}{8^{\frac{1}{3}}} \]

OpenStudy (habbster):

the x in 2.5

OpenStudy (habbster):

thank you a lot

OpenStudy (phi):

and to finish the question the measurement of the edges are 2.5, 15 and 15

OpenStudy (habbster):

yeah i found the answer already ty